OpenStudy (anonymous):

Use the Change of Base Formula to evaluate log8^77. ?

OpenStudy (anonymous):

\[\log_b(A)=\frac{\ln(A)}{\ln(b)}\]

OpenStudy (anonymous):

in your case compute \[\frac{\ln(77)}{\ln(8)}\]

OpenStudy (anonymous):

thank you for your help!